OpenStudy (akgorgeous):

The sum of two consecutive even integers is 234. What is the larger integer?

OpenStudy (jango_in_dtown):

Let the numbers be (n-1) and (n+1). then (n-1)+(n+1)=234 , or, 2n=234 or n=234/2 then n+1 is the larger integer = (234/2) + 1